# Mitigating the Alignment Tax of RLHF

### v3

@@ Abstract
LLMs acquire a wide range of abilities during pre-training, but aligning LLMs under Reinforcement Learning with Human Feedback (RLHF) can lead to forgetting, which is also known as the alignment tax. To empirically verify this hypothesis, we conducted experiments with existing RLHF algorithms using OpenLLaMA-3B, which revealed a pronounced alignment tax in NLP tasks. On the other hand, despite various techniques to mitigate forgetting, they are often at odds with the RLHF performance, leading to a trade-off between reward maximization and forgetting mitigation.
In light of the above pressing issue in aligning LLMs, in this paper we explore model averaging, which interpolates between pre and post RLHF model weights, to achieve a more efficient reward-tax Pareto front. To understand its effectiveness, We offer theoretical insights into model averaging, revealing that it enhances performance Pareto front by increasing feature diversity on the layers where tasks share overlapped feature spaces. Empirical evidence corroborates our analysis by showing the benefits of averaging low-level transformer layers. Building on the analysis and the observation that averaging different layers of the transformer leads to significantly different reward-tax trade-offs, we propose Adaptive Model Averaging (AMA) to adaptively find various combination ratios of model layers. AMA seeks to maximize the alignment reward while incurring minimal alignment tax. Moreover, we validate AMA’s performance across a range of RLHF algorithms over OpenLLaMA-3B and further extend our findings to Mistral-7B.
@@ 1 Introduction
While LLMs exhibit strong abilities among various benchmarks, they still require alignment with human preferences, including the principles of being helpful, honest, and harmless as outlined by (Askell et al., 2021). The goal is to ensure that LLMs are designed to assist users in completing tasks, provide truthful information without deception, and avoid causing harm, whether physical, psychological, or social, to individuals or the environment. The process of aligning LLMs with human preferences often involves the application of Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), as shown in Figure 1. Although RLHF allows LLMs to align with human expectations, prior studies (Askell et al., 2021; OpenAI, 2023; Song et al., 2023) have found that this approach can lead to forgetting in the diverse abilities that the LLMs have already acquired, as illustrated in Figure 1. This phenomenon, also known as the “alignment tax” in the literature, has accumulated substantial attention from both academia and industry (Ouyang et al., 2022; Anthropic, 2023; Askell et al., 2021; Tu et al., 2023; Noukhovitch et al., 2023).
@@
In this paper, we first conduct a comprehensive investigation on alignment tax and develop methods to reduce alignment tax while maintaining the alignment performance (e.g., reward). In particular, we followed the approach presented by (Ouyang et al., 2022) and evaluated alignment tax using multiple NLP benchmarks from common sense QA, such as ARC Easy and Challenge (Clark et al., 2018), Race (Lai et al., 2017), and PIQA (Bisk et al., 2020), reading comprehension benchmarks including SQuAD (Rajpurkar et al., 2018) and DROP (Dua et al., 2019), and translation tasks, including WMT 2014 French to English translation (Bojar et al., 2014) (c.f. Section 2). Our primary focus is on aligning the OpenLLaMA-3B on the helpfulness and harmlessness dataset (Bai et al., 2022) using Rejection Sampling Fine-tuning methods (Dong et al., 2023). Indeed, we observed a substantial alignment tax on these benchmarks consistently, confirming the findings of (Ouyang et al., 2022; Gao et al., 2023). Specifically, as we gained a higher reward during RLHF, indicating better alignment with human preference, the alignment tax also increased simultaneously, clearly inducing a reward-tax trade-off.
@@
We then compare various methods developed in different communities as potential rescues to alleviate the alignment tax. This includes the model averaging method (Wortsman et al., 2022b, a; Lin et al., 2023) from out-of-distribution (OOD) generalization literature, regularization-based techniques from the continual learning literature (Panigrahi et al., 2023; Xuhong et al., 2018; Buzzega et al., 2020; Huang et al., 2021), low-rank adaptation (LoRA) (Huang et al., 2021) from the parameter-efficient fine-tuning literature, as well as the utilization of reward penalty from the reinforcement learning literature (Ziegler et al., 2019; Wu et al., 2021; Ouyang et al., 2022; Yuan et al., 2023). Interestingly, we found that model averaging, which simply interpolates between the weights of models before and after RLHF, achieves the most efficient reward-tax Pareto front. Furthermore, we compared model averaging with Experience Replay (ER), which involves maintaining an additional subset of pre-training datasets during RLHF (Rolnick et al., 2019). Surprisingly, we found that model averaging surpasses ER on two out of three benchmarks (i.e., reading comprehension and translation), despite not being provided with any additional data.
@@
To understand the effectiveness of model averaging, we provide theoretical insights based on the framework of (Lin et al., 2023). In particular, we show that the method can enhance Pareto front by increasing feature diversity on layers where two tasks share similar feature spaces. Empirical evidence also indicates that averaging the low-level layers of Transformers consistently improves both alignment reward and NLP task performance. This aligns with our theoretical insights, as tasks could share similar lower-level features, e.g., better word representation on low-level layers benefits both NLP and alignment tasks.
@@
Additionally, we noticed that averaging different layers of the Transformers unveiled notably distinct patterns of reward-tax trade-off, aligning with our earlier analysis that tasks may exhibit varying overlapping feature spaces in different layers. Motivated by this observation, we propose Adaptive Model Averaging (AMA), which adaptively averages different parts of the models during model averaging. We start by dividing the transformer into $K$ parts and assigning unique averaging ratios for each part, represented as $\alpha_{i}\in[0,1]$ for the $i$th part. AMA aims to maximize alignment reward by optimizing the averaging ratios $(\alpha_{1},\ldots,\alpha_{K})$ while maintaining the overall alignment tax, thus consistently improve the reward-tax Pareto front. To demonstrate the efficiency of AMA, we also contrasted our method with other RLHF techniques, including Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Proximal Policy Optimization (PPO) (Schulman et al., 2017) in Section 5. We further substantiate our findings on Mistral-7B, which further corroborates our empirical findings on OpenLLaMA-3B.
@@
  • We provide theoretical insights into the efficiency of model averaging in enhancing the reward-tax trade-off, demonstrating that both NLP and alignment tasks can benefit from the increased feature diversity from model averaging in the shared feature space.
  • Motivated by our analysis, we introduce Adaptive Model Averaging (AMA), which optimizes the averaging ratios of different model layers to maximize alignment reward. AMA consistently improves the rewards-tax Pareto front across different benchmarks, and it also generalizes well across various RLHF algorithms and different model types, such as OpenLLaMA-3B and Mistral-7B.
The paper is structured as follows: we conduct a systematic investigation of existing methods in Section 2-3. In Section 4, we provide insights into the effectiveness of model averaging. Subsequently, we propose Adaptive Model Averaging in Section 5. We conclude the paper in Section 6.
@@ 2 Experimental Settings
Basic Setting. We chose the OpenLLaMA-3B model (Geng and Liu, 2023) as the base model because of its openly available pre-training dataset, which is essential for investigating the Experience Replay method in Sec. 3.2. Furthermore, we extend the experiments to Mistral-7B in Sec. 5. Following the standard procedure outlined in (Ouyang et al., 2022), we initially conducted instruction tuning, followed by RLHF. Here, $\theta$ represents an LLM with parameters $\theta$, with the pre-trained model denoted as $\theta_{\mathrm{pre}}$. We commenced with instruction fine-tuning for $\theta_{\mathrm{pre}}$ on ShareGPT  [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, which yielded $\theta_{0}$. Subsequently, RLHF was performed on $\theta_{0}$ to obtain $\theta$. Similar to the methodology proposed in (Ouyang et al., 2022), the alignment tax was evaluated by comparing the performance regression of $\theta$ with $\theta_{0}$ across various NLP tasks. The whole procedure and notations are illustrated in Fig. 1.
@@
Figure 2: Illustration of Adaptative Model Averaging (AMA) when $K=3$.
@@ 3 Evaluating Existing Methods
In Figure 12 of Appendix D.1, we visualize the training procedure in terms of the reward-tax trade-off during RLHF. Specifically, we can clearly see that as the RLHF proceeds, the reward begins to increase while the translation and reading comprehension ability continues to drop. Interestingly, we observe that the performance of common sense increases first and then drops. Given that alignment tax is inherently a catastrophic forgetting issue, we then proceed to explore methods to reduce alignment tax. Research focused on reducing forgetting is mainly classified into two main categories, depending on the availability of the pre-training dataset. We also investigate the reward penalty method developed in RL community in Section 3.3.
@@ 3.1 Basic Methods
To explore methods for alleviating alignment tax, we initially examine solutions that do not rely on pre-training datasets. These methods encompass the following:
  • (a)
    Early stopping.
  • (b)
    Regularization towards $\theta_{0}$ in the weight space. We investigate L1 penalty (Xuhong et al., 2018):
    $\max_{\theta}\mathbb{E}_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}[r^{*}(x,a)]+\lambda\|\theta-\theta_{0}\|_{1},$ (2)
    and L2 penalty (Xuhong et al., 2018):
    $\max_{\theta}\mathbb{E}_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}[r^{*}(x,a)]+\lambda\|\theta-\theta_{0}\|_{2},$ (3)
    where $\lambda$ is the penalty weight.
  • (c)
    Low-Rank Adaptation (LoRA) (Hu et al., 2021). It introduces trainable rank decomposition matrices into linear layers to update $\theta-\theta_{0}$ during RLHF.
  • (d)
    Knowledge distillation (Buzzega et al., 2020; Huang et al., 2021). We use $\pi_{\theta_{0}}$ serves as the teacher and $\pi_{\theta}$ as the student, with a penalty imposed as:
@@
  • (e)
    Model Averaging (MA) (Wortsman et al., 2022a, b). This involves simply interpolating between $\theta_{0}$ and $\theta$ to yield the policy $\pi_{(1-\alpha)\theta_{0}+\alpha\theta}$, where $\alpha$ is a hyper-parameter ranging from 0 to 1.
More implementation details are provided in the appendix.
Results. Figure 3 depicts the performance of each aforementioned method. The results demonstrate that these approaches effectively alleviate the alignment tax; however, they also result in a reduction in the RLHF reward, indicating a clear trade-off between reward and alignment tax. Notably, despite its simplicity, the Pareto-front of model averaging supersedes nearly all other methods across various hyper-parameters.
3.2 Experience Replay
During RLHF, it is possible to include a portion of pre-training data (Ouyang et al., 2022), also known as Experience Replay (ER) (Rebuffi et al., ; Shin et al., 2017). However, retaining a subset of the pre-training data entails extra computational costs and implementation intricacies, making it less preferable (Noukhovitch et al., 2023). In this part, we add ER for a comparison with MA. Specifically, we include a small proportion of randomly subsampled pre-training data during the RLHF stage. Here, we denote $\mathcal{D}_{{pre}}$ as the pre-training data distribution, and our objective is to solve the following:
We experiment with different penalty weights $\lambda$ such as $0.25$, $0.5$, $1$, $2$, and $4$. Importantly, we utilize the data proportion as a proxy for setting the penalty weight. For instance, we do not explicitly apply a penalty of 4 when $\lambda=4$; instead, we include 4 times the replay data over the RLHF data in a batch. Refer to the Appendix C for more details.
@@ 3.2 Experience Replay
Figure 4: Comparison of model averaging with Experience Replay.
3.3 Reward Penalty
It is a common practice to impose Kullback–Leibler (KL) penalty on the RL reward in the PPO. Such a penalty can also regularize the policy to stay closer to the initial policy, which in return can reduce the alignment tax. Following (Ziegler et al., 2019; Wu et al., 2021; Ouyang et al., 2022; Yuan et al., 2023), we modify the raw reward function with an additional KL penalty (Ziegler et al., 2019).
where we use $\mathrm{KL}(\pi_{\theta}||\pi_{\theta_{0}})$ to denote $\mathbb{E}_{x}[\mathrm{KL}(\pi_{\theta}(\cdot|x)||\pi_{\theta_{0}}(\cdot|x))]$ for short. We compare vanilla model averaging methods with the reward penalty by considering different KL penalties in $\{0.05,0.1,0.2\}$. The results are shown in Figure 5. We can see that while a larger KL penalty can partially mitigate the forgetting issue, the model averaging is much more effective than the reward penalty in terms of the reward-tax trade-off.
Figure 5: Comparison of model averaging with reward penalty for PPO.
4 Investigating the Mysteries of Model Averaging for Alleviating Alignment Tax

### v4

@@ Abstract
LLMs acquire a wide range of abilities during pre-training, but aligning LLMs under Reinforcement Learning with Human Feedback (RLHF) can lead to forgetting pretrained abilities, which is also known as the alignment tax. To investigate alignment tax, we conducted experiments with existing RLHF algorithms using OpenLLaMA-3B, which revealed a pronounced alignment tax in NLP tasks. Whereas, despite various techniques to mitigate forgetting, they are often at odds with the RLHF performance, leading to a trade-off between alignment performance and forgetting mitigation, leading to an alignment-forgetting trade-off.
In this paper we show that model averaging, which simply interpolates between pre and post RLHF model weights, surprisingly achieves the most strongest alignment-forgetting Pareto front among a wide range of competing methods. To understand its effectiveness, we offer theoretical insights into model averaging, revealing that it enhances performance Pareto front by increasing feature diversity on the layers where tasks share overlapped feature spaces. Empirical evidence corroborates our analysis by showing the benefits of averaging low-level transformer layers. Building on the analysis and the observation that averaging different layers of the transformer leads to significantly different alignment-forgetting trade-offs, we propose Heterogeneous Model Averaging (HMA) to Heterogeneously find various combination ratios of model layers. HMA seeks to maximize the alignment performance while incurring minimal alignment tax. Moreover, we validate HMA’s performance across a range of RLHF algorithms over OpenLLaMA-3B and further extend our findings to Mistral-7B which is evaluated by open-sourced preference model and GPT4. Code available here [^1] https://github.com/avalonstrel/Mitigating-the-Alignment-Tax-of-RLHF.git.
Mitigating the Alignment Tax of RLHF
Yong Lin1thanks: indicates equal contributions, random order. Correspond to <hlinbh@connect.ust.hk>, Hangyu Lin2*, Wei Xiong3*, Shizhe Diao4*, Jianmeng Liu2, Jipeng Zhang2, Rui Pan3, Haoxiang Wang3, Wenbin Hu 2, Hanning Zhang2, Hanze Dong2, Renjie Pi2, Han Zhao3, Nan Jiang3, Heng Ji3, Yuan Yao2, Tong Zhang3 1 Princeton University, Princeton Language and Intelligence 2The Hong Kong University of Science and Technology 3University of Illinois Urbana-Champaign, 4NVIDIA
@@ 1 Introduction
While LLMs exhibit strong abilities among various benchmarks, they still require alignment with human preferences, including the principles of being helpful, honest, and harmless as outlined by (Askell et al., 2021). The goal is to ensure that LLMs are designed to assist users in completing tasks, provide truthful information without deception, and avoid causing harm, whether physical, psychological, or social, to individuals or the environment. The process of aligning LLMs with human preferences often involves the application of Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), as shown in Figure 1. Although RLHF allows LLMs to align with human expectations, prior studies (Askell et al., 2021; OpenAI, 2023; Song et al., 2023) have found that this approach can lead to forgetting in the diverse abilities that the LLMs have already acquired, as illustrated in Figure 1. This phenomenon, also known as the “alignment tax" in the literature, has accumulated substantial attention from both academia and industry (Ouyang et al., 2022; Anthropic, 2023; Askell et al., 2021; Tu et al., 2023; Noukhovitch et al., 2023).
@@
Investigating alignment tax. In this paper, we first conduct a comprehensive investigation on alignment tax and develop methods to reduce alignment tax while maintaining the alignment performance. In particular, we followed the approach presented by (Ouyang et al., 2022) and evaluated alignment tax using multiple NLP benchmarks from common sense QA, such as ARC Easy and Challenge (Clark et al., 2018), Race (Lai et al., 2017), and PIQA (Bisk et al., 2020), reading comprehension benchmarks including SQuAD (Rajpurkar et al., 2018) and DROP (Dua et al., 2019), and translation tasks, including WMT 2014 French to English translation (Bojar et al., 2014) (c.f. Section 3). Our primary focus is on aligning the OpenLLaMA-3B on the helpfulness and harmlessness dataset (Bai et al., 2022) using Rejection Sampling Fine-tuning methods (Dong et al., 2023) (also known as the best-of-$n$ algorithm). In the later part, we extend our experiments to Mistral-7B and Direct Preference Optimization (DPO, Rafailov et al. (2023)). We mainly focus on RSF and DPO since they are popular and nearly all of the latest open-sourced LLMs on the leaderboards are aligned by these two methods [^2] https://tatsu-lab.github.io/alpaca_eval/. Indeed, we observed a substantial alignment tax on these benchmarks consistently, confirming the findings of (Ouyang et al., 2022; Gao et al., 2023). Specifically, as we gained a higher reward during RLHF, indicating better alignment with human preference, the alignment tax also increased simultaneously, clearly inducing a alignment-forgetting trade-off.
@@
Surprising effectiveness of model averaging over. We then compare various methods developed in different communities as potential rescues to alleviate the alignment tax. This includes the model averaging method (Wortsman et al., 2022b, a; Lin et al., 2023) from out-of-distribution (OOD) generalization literature, regularization-based techniques from the continual learning literature (Panigrahi et al., 2023; Xuhong et al., 2018; Buzzega et al., 2020; Huang et al., 2021), low-rank adaptation (LoRA) (Huang et al., 2021) from the parameter-efficient fine-tuning literature, as well as the utilization of reward penalty from the reinforcement learning literature (Ziegler et al., 2019; Wu et al., 2021a; Ouyang et al., 2022; Yuan et al., 2023). Interestingly, we found that model averaging, which simply interpolates between the weights of models before and after RLHF, achieves the most efficient alignment-forgetting Pareto front. In Appendix C.1, we further show and discuss the in-effectiveness of Experience Reply Rebuffi et al. method compared with MA.
@@
Understanding the effectiveness of model averaging. To understand the effectiveness of model averaging, we provide theoretical insights based on the framework of (Lin et al., 2023). In particular, we show that the method can enhance Pareto front by increasing feature diversity on layers where two tasks share similar feature spaces. Empirical evidence also indicates that averaging the low-level layers of Transformers consistently improves both alignment reward and NLP task performance. This aligns with our theoretical insights, as tasks could share similar lower-level features, e.g., better word representation on low-level layers benefits both NLP and alignment tasks.
@@
Heterogeneous model averaging. We noticed that averaging different layers of the Transformers unveiled notably distinct patterns of alignment-forgetting trade-off, aligning with our earlier analysis that tasks may exhibit varying overlapping feature spaces in different layers. Motivated by this observation, we propose Heterogeneous Model Averaging (HMA), which adaptively averages different parts of the models during model averaging. We start by dividing the transformer into $K$ parts and assigning unique averaging ratios for each part, represented as $\alpha_{i}\in[0,1]$ for the $i$th part. HMA aims to maximize alignment reward by optimizing the averaging ratios $(\alpha_{1},\ldots,\alpha_{K})$ while maintaining the overall alignment tax, thus consistently improve the alignment-forgetting Pareto front. To demonstrate the efficiency of HMA, we also contrasted our method with other RLHF techniques, including Direct Preference Optimization (DPO). (Rafailov et al., 2023) We further substantiate our findings on Mistral-7B where evaluations conducted by open sourced perference model and GPT4, which further corroborates our empirical findings on OpenLLaMA-3B.
@@
  • We provide theoretical insights into the efficiency of model averaging in enhancing the alignment-forgetting trade-off, demonstrating that both NLP and alignment tasks can benefit from the increased feature diversity from model averaging in the shared feature space.
  • Motivated by our analysis, we introduce Heterogeneous Model Averaging (HMA), which optimizes the averaging ratios of different model layers to maximize alignment performance. HMA consistently improves the Pareto front across different benchmarks, and it also generalizes well across various RLHF algorithms and different model types, such as OpenLLaMA-3B and Mistral-7B, evaluated by open-sourced preference model and GPT4.
The paper is structured as follows: we conduct a systematic investigation of existing methods in Section 3-4. In Section 5, we provide insights into the effectiveness of model averaging. Subsequently, we propose Heterogeneous Model Averaging in Section 6. We conclude the paper in Section 7.
2 Discussion with existing works.
In this section, we provide comparison of this work with existing works to highlight the novelty of our findings. We defer more comprehensive related works to Appendix A.
Existing works of model averaging for LLMs. Previous research has covered certain aspects of model averaging. Ramé et al. (2024) demonstrate the utilization of model averaging to construct a more resilient reward model for reinforcement learning with human feedback (RLHF). In a similar vein, Rame et al. (2024) employ model averaging to merge policy models trained for distinct objectives, facilitating multi-objective RLHF. Sanyal et al. (2023) introduce the integration of moving averaging to enhance pre-training. However, none of these studies investigate the alignment tax, and their findings are independent of our research.
Existing works on finding adaptive combinations for model merging. Previous studies Yang et al. (2023); Akiba et al. (2024) have also discussed the idea of dynamically assigning different weights to different layers when merging models, aiming to maximize performance on a specific task (e.g., $\mathcal{T}_{i}$). These approaches assume access to the task-specific data $\mathcal{T}_{i}$. However, considering the nature of alleviating alignment tax, which aims to mitigate forgetting across a extremely wide range of tasks ($\mathcal{T}_{j_{1}}...\mathcal{T}_{j_{K}}$), these methods fail to effectively optimize performance for multiple tasks simultaneously. In the Appendix E.4, we demonstrate that using the method proposed by Yang et al. (2023), which optimizes for a single task, does not effectively address forgetting on the other tasks. Furthermore, our work is the first to provide an explanation for the surprising effectiveness of model averaging in alleviating forgetting, as well why we should assign heterogeneous combination ratios.
Existing works on the forgetting of language models. Most research on forgetting in language models focuses on sequentially pre-training Chen et al. (2023); Gong et al. (2022); Jin et al. (2021); Qin et al. (2022); Liu et al. (2021) or fine-tuning tasks Sun et al. (2019); Razdaibiedina et al. (2023); Wu et al. (2021b); Zhang et al. (2022); Madotto et al. (2020), e.g., sequentially training on task $\mathcal{T}_{i}$ and then task $\mathcal{T}_{j}$. They evaluate forgetting by measuring the model’s performance on a task (e.g., task $\mathcal{T}_{i}$) after training it on another task (e.g., task $\mathcal{T}_{j}$). However, these methods have not explored the effectiveness of model averaging. In our case, we demonstrate the significant power of model averaging which outperform a wide range of existing methods. Furthermore, existing works assume that the data size of each task is comparable (i.e., the dataset size of $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ is similar), allowing for a subset (e.g., 10%) of old task data replay, which is shown to effective alleviate the forgetting without excessive computation overhead in their settings. However, in our alignment tax situation, we aim to preserve a wide range of abilities gained during pre-training, which is challenging since pre-training datasets are often not publicly available. In Appendix C.1, we show that even when we have access to the pre-training data and replay a subset up to four times larger than the RLHF data (which costs significant computation overhead), experience replay still under-performs model averaging in two out of three benchmarks. This is likely due to the vast size of the pre-training data, where the subset only covers a small fraction of it (e.g., only covers ~$0.01\%$ of the pre-training data). So replay methods are less practical for alleviating alignment tax.
@@ 3 Experimental Settings
Basic Setting. We chose the OpenLLaMA-3B model (Geng and Liu, 2023) because (1) it is computational friendly compared with 7B models (2) it has openly available pre-training dataset, which is convenient to investigate Experience Replay in Appendix. C.1. Furthermore, we extend the experiments to Mistral-7B in Sec. 6. Following the standard procedure outlined in (Ouyang et al., 2022), we initially conducted instruction tuning, followed by RLHF. Here, $\theta$ represents an LLM with parameters $\theta$, with the pre-trained model denoted as $\theta_{\mathrm{pre}}$. We commenced with instruction fine-tuning for $\theta_{\mathrm{pre}}$ on ShareGPT  [^3] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, which yielded $\theta_{0}$. Subsequently, RLHF was performed on $\theta_{0}$ to obtain $\theta$. Similar to the methodology proposed in (Ouyang et al., 2022), the alignment tax was evaluated by comparing the performance regression of $\theta$ with $\theta_{0}$ across various NLP tasks. The whole procedure and notations are illustrated in Fig. 1.
@@
Figure 2: Illustration of Heterogeneous Model Averaging (HMA) when $K=3$.
@@ 4 Evaluating Existing Methods
In Figure 12 of Appendix E.1, we visualize the training procedure in terms of the alignment-forgetting trade-off during RLHF. Specifically, we can clearly see that as the RLHF proceeds, the reward begins to increase while the translation and reading comprehension ability continues to drop. Interestingly, we observe that the performance of common sense increases first and then drops. Given that alignment tax is inherently a catastrophic forgetting issue, we then proceed to explore methods to reduce alignment tax. Research focused on reducing forgetting is mainly classified into two main categories, depending on the availability of the pre-training dataset. We also investigate the reward penalty method developed in RL community in Appendix C.2.
@@ 4.1 Basic Methods
To explore methods for alleviating alignment tax, we initially examine solutions that do not rely on pre-training datasets. These methods encompass the following:(a) Early stopping. (b) Regularization towards $\theta_{0}$ in the weight space as follows:
$\max_{\theta}\mathbb{E}_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}[r^{*}(x,a)]+\lambda\|\theta-\theta_{0}\|_{\alpha},$ (2)
where we use $\alpha=1,2$ which corresponds to the L1 and L2 (Xuhong et al., 2018) penalties, respectively. (c) Low-Rank Adaptation (LoRA) (Hu et al., 2021). It introduces trainable rank decomposition matrices into linear layers to update $\theta-\theta_{0}$ during RLHF. (d) Knowledge distillation (Buzzega et al., 2020; Huang et al., 2021). We use $\pi_{\theta_{0}}$ serves as the teacher and $\pi_{\theta}$ as the student, with a penalty imposed as:
@@
(e) Model Averaging (MA) (Wortsman et al., 2022a, b). This involves simply interpolating between $\theta_{0}$ and $\theta$ to yield the policy $\pi_{(1-\alpha)\theta_{0}+\alpha\theta}$, where $\alpha$ is a hyper-parameter ranging from 0 to 1. (f) Stochastic Moving Averaging (SMA) Noukhovitch et al. (2024). More implementation details are provided in the appendix.
Results. Figure 3 depicts the performance of each aforementioned method. The results demonstrate that these approaches effectively alleviate the alignment tax; however, they also result in a reduction in the RLHF reward, indicating a clear trade-off between reward and alignment tax. Notably, despite its simplicity, the Pareto-front of model averaging supersedes nearly all other methods across various hyper-parameters. In Appendix C.1 and C.2, we compared model averaging with Experience Replay (ER) and KL reward penalty methods for Proximal policy optimization Schulman et al. (2017) algorithms, the conclusions are similar.
@@
5 Unravelling the Mysteries of Model Averaging for Alleviating Alignment Tax